# Nightjar: Dynamic Adaptive Speculative Decoding for Large Language Models Serving

### v1

@@ 3.1 Experiment setup
Fig. 5: Method Comparison at Low Request Rate.
Fig. 6: Method Comparison at High Request Rate.
@@
Fig. 7: Distribution of Input and Output Lengths Across Three Datasets.
@@
Settings. We select DeepSeek-R1-Distill-Qwen-7B [deepseek2024deepseekr1] and vicuna-13b [vicuna13b] as target models and match them with DeepSeek-R1-DRAFT-Qwen2.5-0.5B [alamios_DeepSeek_R1_DRAFT_2024] and vicuna-68m [double7_vicuna68m] as draft models respectively. The model pair for the 7B target model is operated on a single RTX 4090 (24GB) GPU, and the model pair for the 13B target model operates on an A100 (40GB) GPU. Nightjar is implemented by extending the vLLM (version 0.8.2) [kwon2023efficient].
@@
Workloads. We selected real-world production traces from the AzureLLMInferenceDataset [AzurePublicDataset], and chose a segment with dynamic request rates as shown in Fig. 9. For online chatting, we utilize datasets from ShareGPT [anon8231489123_ShareGPT_V3], Alpaca [tatsu-lab_alpaca]. For other tasks, we utilize SpecBench [xia-etal-2024-unlocking], which is composed of 480 randomly selected instances from six widely used databases. For each dataset, we use 480 instances. The input and output length distribution is presented in Fig. 7.
[…]
Fig. 8: Throughput comparison for static request rate on SpecBench dataset. (Left) Comparison between SOTA algorithms. (Right) Comparison among different speculative length $\gamma$.
3.2 Static request load
@@
Nightjar demonstrates robust performance across varying load conditions by dynamically adapting its speculation strategy. In the detailed breakdown by request rate on ShareGPT dataset, under low request rates (Fig. 5) where Speculative Decoding (SD) outperforms standard decoding, Nightjar maximizes efficiency by selecting the optimal draft length. Conversely, in high request rate scenarios (Fig. 6) where the overhead of SD renders it inferior to the baseline, Nightjar adaptively disables speculation to avoid performance degradation, thereby consistently achieving superior results in both regimes. Note that data for $\gamma=5$ at 20 QPS is excluded from Fig. 6 owing to GPU memory limitations. These advantages are similarly observed on SpecBench dataset (Fig. 8); specifically, Nightjar outperforms baselines at high concurrency (e.g., 25 QPS in Fig. 8) and matches the throughput of the optimal fixed speculative length ($\gamma=4$) in lower traffic conditions (Fig. 8). This consistent superiority stems from Nightjar’s bandit policy, which dynamically optimizes the speculation length to balance verification costs and generation speed, ensuring optimal efficiency whether the request load is low or high.
@@ 3.3 Dynamic request load
Fig. 9: Request Rate Trace.
Fig. 10: Throughput Trace.
For 7B model under dynamic request rate (Fig. 9), as shown in Fig. 10, Nightjar surges ahead during low request load periods, achieving the highest peak throughput ( 3800 tokens/s) and maintaining a consistent lead over rivals during high load phases. DSD’s predictions are less accurate during low request load, leading to throughput degradation, but perform better under high request load, while TETRIS exhibits performance drops under high request load due to its high overhead in large batch size and static BanditSpec is not suitable for dynamic batch size.
Table 2: Throughput (tokens/s) Comparison for 7B and 13B Models
@@
Table 3: Mean End-to-End Latency (ms) Comparison for 7B and 13B Models
@@
As shown in Table 2, Nightjar consistently achieves the highest throughput, demonstrating superior scalability over both static and dynamic baselines. While Standard Speculative Decoding (SD) provides moderate gains, Nightjar significantly amplifies these benefits. Notably, on the 13B model, Nightjar achieves a remarkable 64.6% improvement over vanilla decoding (w/o SD) on the Alpaca dataset. More importantly, Nightjar shows exceptional robustness in complex scenarios. Compared to the advanced inference system TETRIS, Nightjar maintains a consistent lead, achieving a 12.1% throughput gain on the 7B Alpaca dataset. Furthermore, on the 13B SpecBench dataset, Nightjar outperforms the competing DSD method by a substantial 53.4% and BanditSpec by 39.1%. Even against the optimized fixed-length SD baseline, Nightjar maintains a clear advantage, proving that its cost-aware policy effectively identifies optimal lengths.
@@
Regarding end-to-end latency, Table 3 further validates Nightjar’s efficiency. For the 7B model, Nightjar reduces latency by 20.2% compared to SD on the Alpaca dataset. Crucially, Nightjar demonstrates superior lightweight scheduling compared to TETRIS. On the 7B Alpaca dataset, Nightjar reduces latency by 24.4% compared to TETRIS, indicating that Nightjar effectively mitigates the complex sampling overheads that hinder TETRIS. On the 13B model, the advantage remains significant against other dynamic baselines: Nightjar reduces latency by 34.3% compared to DSD on SpecBench. This confirms that Nightjar’s bandit-based mechanism avoids the ”decision deadlock” and high penalty costs that hamper DSD in high-load, compute-bound environments.
Fig. 11: Performance Comparison of Context-Aware Bandit Methods.
3.4 Ablation Studies
@@
Fig. 11 additionally compares various context-aware multi-armed bandit methods (batch size as context) and the basic ADA-BINGREEDY approach. Our method consistently achieves optimal performance across different QPS levels because LinUCB and linearTS assume linear relationships between rewards and contexts, which don’t hold for LLM speculative serving scenarios. Throughput depends not only on hardware but also on acceptance rates and model types, making linear models unsuitable for this complex scenario.

### v2

@@ 24 $t\leftarrow t+1$
4.3 Theoretical Analysis
In this section, we provide a rigorous theoretical analysis for our proposed Nightjar algorithm. We begin by presenting our main theoretical result, Theorem 1, which establishes a sublinear regret bound for the algorithm under the dynamic serving environment. The proof relies on decomposing the regret into exploration, estimation, and switching cost components.
For the analysis, we assume the goodput reward $r_{t}$ is bounded in $0,g_{\max}]$ and the switching cost is bounded by $C_{\text{switch}}$.
4.3.1 Regret of Nightjar
The core element of our proposed framework is the adaptive bin-greedy strategy. The following theorem presents the regret upper bound of Nightjar.
Theorem 1.
For any time horizon $T$ and batch size context $B$, assuming the switching cost is bounded, there exist positive constants such that with probability at least $1-\delta$, the cumulative regret of the proposed Nightjar algorithm satisfies:
$R(T)\leq\mathcal{O}\left(g_{\max}\sqrt{T\log T}+C_{\text{switch}}\sqrt{T}\right)=\tilde{\mathcal{O}}(\sqrt{T}).$ (4)
Consequently, the average regret satisfies $\lim_{T\to\infty}R(T)/T=0$.
Remark 1.
Theorem 1 provides a high-probability theoretical guarantee. Unlike expected regret bounds, this ensures that the algorithm performs efficiently in widely varying single runs, which is critical for system stability in production environments.
4.3.2 Proof of Theorem 1
To prove Theorem 1, we introduce three supporting lemmas handling estimation confidence, exploration frequency, and switching frequency.
Lemma 1 (Confidence Bound).
For any arm $\gamma$ and time $t$, let $N_{t}(\gamma)$ be the number of times arm $\gamma$ has been pulled. With probability at least $1-\delta/2$, the deviation between the empirical mean goodput $\tilde{g}_{t,\gamma}$ and the true mean $\mu_{\gamma}$ is bounded by the confidence radius $\beta_{t}(\gamma)$:
$|\tilde{g}_{t,\gamma}-\mu_{\gamma}|\leq\beta_{t}(\gamma)\coloneqq\sqrt{\frac{g_{\max}^{2}\log(2T/\delta)}{2N_{t}(\gamma)}}.$ (5)
Lemma 2 (Bounding Exploration).
Under the adaptive binning mechanism, the exploration probability in the $b$-th bin is set exactly to $p_{b}=1/\sqrt{b}$. The cumulative number of exploration steps up to time $T$ is bounded by:
$N_{\text{explr}}(T)=\sum_{t=1}^{T}\mathbb{I}(\text{exploration}_{t})\leq\mathcal{O}(\sqrt{T}).$ (6)
Lemma 3 (Bounding Switching Cost).
The total number of speculative length switches, denoted as $S(T)$, is bounded by the sum of exploration switches and exploitation updates. Given the block-bin structure where block size grows geometrically, $S(T)$ satisfies:
$S(T)\leq\mathcal{O}(\sqrt{T}).$ (7)
Indeed, switches primarily occur during exploration steps (bounded by Lemma 2) or at bin boundaries; since bin sizes grow with $\sqrt{H_{B}}$, the frequency of boundary updates decays, which yields $S(T)\leq\mathcal{O}(\sqrt{T})$.
Based on Lemmas 1, 2, and 3, we derive Theorem 1.
Proof of Theorem 1.
The cumulative regret $R(T)$ can be decomposed into three parts: regret from exploration rounds ($R_{\text{explr}}$), regret from suboptimal choices during exploitation due to estimation error ($R_{\text{est}}$), and the cumulative switching cost ($C_{\text{switch}}\cdot S(T)$).
$R(T)$ $=R_{\text{explr}}+R_{\text{est}}+\text{SwitchingCost}(T)$
$\leq\underbrace{N_{\text{explr}}(T)\cdot g_{\max}}_{(a)}+\underbrace{\sum_{t\in\mathcal{T}_{\text{exploit}}}(\mu^{*}-\mu_{\gamma_{t}})}_{(b)}+\underbrace{S(T)\cdot C_{\text{switch}}}_{(c)}$
Term (a): By Lemma 2, the exploration regret is bounded by $g_{\max}\cdot\mathcal{O}(\sqrt{T})$.
Term (c): By Lemma 3, the switching cost is bounded by $C_{\text{switch}}\cdot\mathcal{O}(\sqrt{T})$.
Term (b): For exploitation steps, if a suboptimal arm $\gamma_{t}$ is chosen, it implies the empirical estimate $\tilde{g}_{t,\gamma_{t}}$ exceeded the optimal $\tilde{g}_{t,\gamma^{*}}$. Using the confidence bound from Lemma 1, the instantaneous regret is bounded by $2\beta_{t}(\gamma_{t})$. Summing over $T$, standard bandit analysis (deriving from the summation of inverse square roots) yields:
$\sum_{t=1}^{T}2\sqrt{\frac{g_{\max}^{2}\log(2T/\delta)}{2N_{t}(\gamma_{t})}}\leq\mathcal{O}\left(g_{\max}\sqrt{T\log T}\right).$ (8)
Combining these terms:
$R(T)$ $\leq\mathcal{O}(g_{\max}\sqrt{T})+\mathcal{O}(g_{\max}\sqrt{T\log T})+\mathcal{O}(C_{\text{switch}}\sqrt{T})$
$=\mathcal{O}\left(\sqrt{T\log T}\right).$ (9)
This completes the proof. ∎
5 Elastic Memory Management for Model-Based Draft Models
Figure 6: (a) vLLM mitigates memory fragmentation via KV cache virtualization but spatially isolates activation values from the KV cache. (b) Nightjar logically decouples KV cache management from model weights while consolidating them into a unified physical memory pool to maximize utilization.
Under high system contention, characterized by a backlog in the request queue, the scarcity of KV cache memory often necessitates frequent swap operations or the rejection of active requests. In such saturation regimes, speculative decoding becomes counter-productive due to scheduling overheads. To address this, we propose a dynamic offloading mechanism for model-based draft models. As illustrated in Figure 6, we offload the draft model weights—analogous to LoRA adapter management—to the host memory, thereby reclaiming GPU memory to expand the capacity of the original KV cache. Let $M_{\text{orig}}$ denote the baseline KV cache capacity; the expanded capacity is defined as $M_{\text{scale}}=M_{\text{orig}}+S_{\text{draft}}$, where $S_{\text{draft}}$ represents the memory footprint of the draft model.
5.1 Trigger Conditions for Dynamic Adjustment
The system employs an elasticity policy based on real-time metrics of KV cache occupancy and queue depth. A KV cache expansion event is triggered when the following conditions converge: (1) Speculative decoding is disabled; (2) The number of available memory blocks, denoted as $N_{\text{free}}$, falls below a critical threshold $\tau_{\text{low}}$ ($N_{\text{free}}<\tau_{\text{low}}$); (3) This resource scarcity persists for a temporal window of $T_{\text{persist}}$ scheduling steps.
Conversely, a KV cache contraction is initiated to restore the draft model when: (1) The waiting queue is empty ($|\mathcal{Q}_{\text{wait}}|=0$); (2) Sufficient memory exists to accommodate both the draft model and a safety buffer ($N_{\text{free}}>S_{\text{draft}}+\tau_{\text{low}}$). This hysteresis design ensures that speculative decoding remains disabled during high-load phases (where $N_{\text{free}}<\tau_{\text{low}}$), thereby preventing resource thrashing.
5.2 Mechanism of KV Cache Expansion
Upon triggering expansion, the system dynamically reallocates GPU memory. The Model Executor allocates a contiguous memory segment of size $S_{\text{draft}}$ to the KV cache pool. Subsequently, the Block Manager updates its internal state: (1) The set of allocatable physical block indices is expanded to include the new range $K_{\text{curr}},K_{\text{new}})$; (2) Reference counters for these newly instantiated blocks are initialized to zero; (3) The new block indices are appended to the free block queue. Crucially, this process is implemented via in-place expansion, eliminating the need for massive memory copying and minimizing reconfiguration overhead.
5.3 KV Cache Contraction and Logical Remapping
Figure 7: Overview of the KV Cache Block Contraction and Logical Table Remapping process. The system ensures data consistency during physical memory compaction through topological sorting and vectorized data migration.
Contraction is the most critical phase, requiring the compaction of the physical memory layout without compromising active contexts. As depicted in Figure 7, the process involves four sequential steps:
Step 1: Identification of Eviction Candidates. The system scans the page tables of all active sequences to identify the set of physical blocks $\mathcal{B}_{\text{evict}}$ located in the memory region to be reclaimed (i.e., blocks with $\text{ID}\geq K_{\text{new}}$). These blocks must be migrated to the preserved memory region $0,K_{\text{new}})$.
Step 2: Construction of Migration Plan. We verify available slots in the free list with IDs strictly less than $K_{\text{new}}$. A mapping $\mathcal{M}:b_{\text{old}}\rightarrow b_{\text{new}}$ is constructed. To preserve logical dependencies (specifically, the pointers to predecessor blocks), we apply a topological sort to the migration set. This ensures that blocks are migrated in an order that maintains the integrity of the dependency chain.
Step 3: Metadata Update and Remapping. For each block requiring migration, the system performs atomic metadata updates: (a) A new physical block object is instantiated, inheriting the token sequence and size of the original; (b) Predecessor pointers are updated—if a predecessor block is also part of the migration set, the pointer is redirected to its new location; otherwise, the original reference is maintained; (c) The logical Block Table is updated to map the sequence to the new physical block ID; (d) The reference counter for the new block is incremented, and it is marked as allocated. Finally, the old physical block is released, ensuring the logical sequence remains consistent with the physical memory state.
Step 4: Vectorized Data Migration. Following the metadata update, the physical data transfer is executed. To handle potential circular dependencies in the mapping (e.g., $A\to B$ and $B\to A$), the topological sort resolves execution order, isolating any simplified cyclic swaps. The migration is powered by a custom high-performance kernel (implemented in Triton) utilizing a tile-based parallelization strategy. Each thread block manages the transfer of a complete KV cache block via vectorized load/store instructions. This approach provides robust performance even on hardware lacking advanced virtual memory management (VMM) features like cuMemMap. By decoupling memory reclamation from model execution, we eliminate deadlock risks associated with ”hold-and-wait” patterns.
Step 5: Allocator State Finalization. Post-migration, the block allocator trims its index set, removing IDs $\geq K_{\text{new}}$ and clearing associated reference counts, effectively finalizing the contraction.
5.4 Non-blocking Asynchronous Migration
To minimize latency, we leverage the fact that speculative decoding is inactive during the unloading phase. Consequently, draft model weights are evicted from the GPU immediately, with a backup retained in host memory. Conversely, when VRAM availability indicates a low request rate, the draft model is reloaded to resume accelerated generation. This reloading process utilizes asynchronous CUDA streams with non-blocking Direct Memory Access (DMA). This allows data transfer to overlap with ongoing computation, ensuring that the background management of model weights incurs negligible impact on the serving latency of the primary model.
@@ 6.1 Experiment setup
Figure 8: Distribution of Input and Output Lengths Across Three Datasets.
@@
Settings. We select DeepSeek-R1-Distill-Qwen-7B [7] and Vicuna-13b [21] as target models and match them with DeepSeek-R1-DRAFT-Qwen2.5-0.5B [2] and vicuna-68m [8] as draft models respectively. The model pair for the 7B target model is operated on a single RTX 4090 (24GB) GPU, and the model pair for the 13B target model operates on an A100 (40GB) GPU. Nightjar is implemented by extending the vLLM (version 0.8.2) [13].
@@
Workloads. To simulate a dynamic user environment, we generate request arrivals using a Poisson distribution. We also selected real-world production traces from the AzureLLMInferenceDataset [4], and chose a segment with dynamic request rates as shown in Figure 10. For online chatting, we utilize datasets from ShareGPT [3], Alpaca [14]. For other tasks, we utilize SpecBench [28], which is composed of 480 randomly selected instances from six widely used databases. For each dataset, we use 480 instances. The input and output length distribution is presented in Figure 8.
Baselines. We evaluate the total token throughput (tokens/s) and mean end-to-end latency of Nightjar against baseline methods:
  • Standard Speculative Decoding (SD): Implemented on top of vLLM, this baseline employs a vanilla chain-style draft generation (e.g., $\gamma=3$), where the draft model generates a sequence of candidate tokens followed by parallel verification from the target model.
  • Vanilla Decoding (w/o SD): This baseline does not use speculative decoding, where the target model directly generates the response tokens without using the draft model.
  • Dynamic Speculative Decoding (DSD) [19]: This method utilizes a linear regression model to predict execution latency and estimates the expected acceptance rate based on historical statistics. It dynamically selects the speculative length that maximizes the system goodput, defined as the ratio of the expected number of accepted tokens (accepted length +1+1) to the execution time.
  • BanditSpec [10]: This approach treats the selection of speculative length as a Multi-Armed Bandit (MAB) problem, leveraging a variant of the Upper Confidence Bound (UCB) algorithm for optimization. However, it does not explicitly incorporate batch size as a contextual feature in its decision-making process.
  • TETRIS [27]: Given a fixed speculative length (similar to standard SD), TETRIS keeps the draft generation process intact but re-engineers the verification phase. Within a predefined computational budget, it prioritizes the verification of tokens with the highest predictive entropy across a batch to improve verification efficiency.
Our result is averaged over 5 independent runs.
Figure 9: Method Comparison at Low Request Rate and High Request Rate for 7B model.
@@
Nightjar demonstrates robust performance across varying load conditions by dynamically adapting its speculation strategy. As shown in Figure 9, in the detailed breakdown by request rate on ShareGPT dataset, under low request rates (Request Rate ¡ 4) where speculative decoding outperforms standard decoding, Nightjar maximizes efficiency by selecting the optimal draft length. Conversely, in high request rate scenarios where the overhead of SD renders it inferior to the baseline, Nightjar adaptively disables speculation to avoid performance degradation, thereby consistently achieving superior results in both regimes. Note that data for $\gamma=5$ at 20 QPS is excluded from Figure 9 owing to GPU memory limitations.
@@ 6.2 Dynamic request load
Figure 10: Request Rate Trace.
Figure 11: Throughput Trace for 7B model under dynamic request rate.
For 7B model under dynamic request rate (Figure 10), as shown in Figure 11, Nightjar surges ahead during low request load periods, achieving the highest peak throughput ( 3800 tokens/s) and maintaining a consistent lead over rivals during high load phases. DSD’s predictions are less accurate during low request load, leading to throughput degradation, but perform better under high request load, while TETRIS exhibits performance drops under high request load due to its high overhead in large batch size and static BanditSpec is not suitable for dynamic batch size.
TABLE IV: Throughput (tokens/s) Comparison for 7B and 13B Models
@@
TABLE V: Mean End-to-End Latency (ms) Comparison for 7B and 13B Models
@@
As shown in Table IV, Nightjar consistently achieves the highest throughput, demonstrating superior scalability over both static and dynamic baselines. While SD provides moderate gains, Nightjar significantly amplifies these benefits. Notably, on the 13B model, Nightjar achieves a remarkable 64.6% improvement over vanilla decoding (w/o SD) on the Alpaca dataset. More importantly, Nightjar shows exceptional robustness in complex scenarios. Compared to the advanced inference system TETRIS, Nightjar maintains a consistent lead, achieving a 12.1% throughput gain on the 7B Alpaca dataset. Furthermore, on the 13B SpecBench dataset, Nightjar outperforms the competing DSD method by a substantial 53.4% and BanditSpec by 39.1%. Even against the optimized fixed-length SD baseline, Nightjar maintains a clear advantage, proving that its cost-aware policy effectively identifies optimal lengths.
@@
Regarding end-to-end latency, Table V further validates Nightjar’s efficiency. For the 7B model, Nightjar reduces latency by 20.2% compared to SD on the Alpaca dataset. Crucially, Nightjar demonstrates superior lightweight scheduling compared to TETRIS. On the 7B Alpaca dataset, Nightjar reduces latency by 24.4% compared to TETRIS, indicating that Nightjar effectively mitigates the complex sampling overheads that hinder TETRIS. On the 13B model, the advantage remains significant against other dynamic baselines: Nightjar reduces latency by 34.3% compared to DSD on SpecBench. This confirms that Nightjar’s bandit-based mechanism avoids the ”decision deadlock” and high penalty costs that hamper DSD in high-load, compute-bound environments.
6.3 Ablation and Analysis
6.3.1 Context-Aware Bandit Methods
(a) Alpaca
(b) ShareGPT
(c) SpecBench
Figure 12: Context-aware bandit method comparison on Alpaca, ShareGPT and SpecBench datasets.
@@ 6.3.1 Context-Aware Bandit Methods
We conduct ablation experiments on 7B model. Figure 12 additionally compares various context-aware multi-armed bandit methods (batch size as context) and the basic ADA-BINGREEDY approach. Our method consistently achieves optimal performance across different QPS levels because LinUCB and linearTS assume linear relationships between rewards and contexts, which don’t hold for LLM speculative serving scenarios. Throughput depends not only on hardware but also on acceptance rates and model types, making linear models unsuitable for this complex scenario.
Figure 13: Performance Comparison of with offload and without offload
6.3.2 Efficacy of Dynamic Model Offloading
The impact of our dynamic loading and unloading mechanism is assessed by comparing Nightjar-Offload with the standard Nightjar implementation. Results in Figure 13 demonstrate that offloading becomes increasingly beneficial as the system enters the compute-bound regime at high request rates, reaching a peak throughput of 6315.9 tok/s at 35 req/s—a substantial improvement over the 5982.6 tok/s achieved without offloading. This confirms that liberating GPU memory by unloading draft models effectively expands the KV cache capacity for the target model, directly translating to higher service concurrency.
Figure 14: Performance Comparison of with different thresholds
6.3.3 Sensitivity to Trigger Thresholds
We further investigate the sensitivity of system performance to the free KV cache threshold $N_{low}$, which dictates the timing for dynamic adjustments. Figure 14 illustrates that the system reaches an optimal throughput plateau of approximately 6110 tok/s at a 10% free cache threshold, which significantly exceeds the fixed baseline performance of 5768.12 tok/s. This result indicates that a 10% buffer provides the ideal balance between sustaining speculative decoding and preventing memory exhaustion, ensuring robust QoS under heavy traffic conditions.
6.3.4 Compare with different speculative length
Figure 15: Throughput comparison for static request rate on SpecBench dataset. Comparison among different speculative length $\gamma$ for 13B model.
As shown in Figure 15, when compared against the best-performing fixed speculative lengths (such as $\gamma=2$ or $\gamma=4$ at their peak efficiency) for 13B model, Nightjar consistently maintains a 1% to 16% performance advantage. This demonstrates exceptional robustness, as Nightjar never falls behind the optimal static configuration.
6.3.5 Scalability Analysis on Multi-GPU Settings
Figure 16: Performance Comparison of 30B model
In this experiment, we deploy a 30B-scale model across two NVIDIA L20 (48GB) GPUs using tensor parallelism. Specifically, we use Qwen2.5-32B-Instruct as the target model and Qwen2.5-0.5B-Instruct as the draft model. As shown in Figure 16, by comparing Nightjar against other baselines, we observe that Nightjar maintains its performance advantage even when scaling to larger models distributed across multiple GPUs.
6.3.6 Overhead of Elastic Memory Operations
TABLE VI: Overhead of Elastic Memory Operations for 30B model on 2 L20 GPUs.
Operation Latency
KV Cache Block Expansion 143.9 ms
KV Cache Block Contraction (Triton-accelerated) 11.9 ms
[…]
7 Related Work
7.1 Speculative Decoding in LLM Serving
While various approaches have sought to optimize speculative decoding, they often fail to address the dynamic and unpredictable nature of real-world LLM serving loads. Initial efforts focused on single-request optimization [26, 31, 32, 6, 11] or static offline batches [25, 10], employing predictive algorithms [31] or confidence-based methods [12, 26, 27], but their static nature makes them fundamentally ill-suited for the fluctuating request rates in live environments. More recent works have explored dynamic speculation length adjustment in serving systems [25, 19, 12, 18], yet they exhibit significant limitations. For instance, SpecServe [12] and TETRIS [27] rely on draft confidence scores, but this can lead to poor performance with large batch sizes and still requires generating at least one token. DSD [19] optimizes length via linear goodput modeling, but its reliance on past average acceptance rates introduces a critical vulnerability. This dependency on historical data, found in methods like DSD [19] and SpecServe [12], can lead to a “decision deadlock” where disabling speculative decoding permanently halts the collection of new acceptance rate data, locking the system in a suboptimal state. Our exploration-exploitation approach is designed to mitigate this risk. Even the work of BanditSpec [10], which first introduced a Multi-Armed Bandit (MAB) paradigm, is constrained by its static design, rendering it incompatible with the continuous batching [13] of modern serving systems. Furthermore, to the best of our knowledge, existing works largely overlook the “switching cost” incurred when re-enabling speculative decoding from a length of 0, which is the overhead of KV cache reconstruction for the draft model.
7.2 Unified Memory Management for Weights and KV Cache in LLM Serving
Oneiros [17] targets multi-model deployment by introducing a dynamic memory remapping engine that redirects GPU memory reserved for model parameters to the KV cache. Unlike Nightjar, which completely unloads draft model weights when they are not in use , Oneiros operates on active models. By performing remapping at layer granularity, it enables models to remain in a serving state while utilizing asynchronous transfers to fetch required parameters in real-time before execution. This approach effectively hides latency by overlapping computation with parameter transfer. eLLM [29] focuses on intra-model efficiency, enabling the KV cache and activation tensors to ”borrow” space from a shared physical memory pool, effectively breaking the traditional isolation between these resources. FineServe [5] provides a unified management framework for both weights and KV caches across multiple models, specifically addressing the challenges of serving models with heterogeneous quantization precisions.